OpenStudy (anonymous):

Ten is the geometric mean of 4 and what number? Answer 2

OpenStudy (asnaseer):

you should be able to use the fact stated in your previous question to work it out.

OpenStudy (anonymous):

how do i do this one because it doesnt give me the two numbers. is it 25?

OpenStudy (asnaseer):

yes 25 is the answer

OpenStudy (anonymous):

thanks a lot again

OpenStudy (asnaseer):

in cases like this you should create a symbol for the unknown quantity - call the unknown number x. then you can write:\[10=\sqrt{4x}\]

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

Would proportions help? 4/10 = 10/x or 4x = 100